OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (anonymous):

ok so i know that it will look like (x-2)(x+4)(x+(1+3i))(x-(1+3i)) but i don't know how to factor the last half

OpenStudy (anonymous):

(x^2+2x-8)(x+(1+3i))(x-(1+3i))

OpenStudy (anonymous):

@precal

OpenStudy (precal):

you need to expand it, do you know how to work with complex numbers

OpenStudy (anonymous):

sorta, i know how to do it but i never get the answer right

OpenStudy (precal):

no your complex solutions are 1+3i and 1 - 3i

OpenStudy (precal):

those are your conjugates

OpenStudy (anonymous):

well, now i know why i always get it wrong...

OpenStudy (precal):

(1+3i)(1-3i) 1^2+3i-3i-9i^2 1-9i^2 1-9(-1) 1+9 10 I think it is x^2-10

OpenStudy (precal):

let me think about this........

OpenStudy (precal):

not x^2-10

OpenStudy (anonymous):

f(x) = x4 - 2x2 + 36x - 80 f(x) = x4 - 3x3 + 6x2 - 18x + 80 f(x) = x4 - 9x2 + 36x - 80 f(x) = x4 - 3x3 - 6x2 + 18x - 80 these are the possible answers

OpenStudy (precal):

that doesn't help

OpenStudy (precal):

x^2+10=0 using quadratic formula a=1 b=0 c=10

OpenStudy (precal):

x=-0+square root of (0^2-4(1)(10) ) all over 2(1) x=square root of -40 all over 2 \[\frac{ \sqrt{-40} }{ 2 }\]

OpenStudy (precal):

Let me go look at some of my old classnotes, be right back

OpenStudy (anonymous):

ok, thank you

OpenStudy (precal):

oh wow this one is in my book

OpenStudy (precal):

[x-(1+3i)][x-(1-3i)] they expanded this out [x-1-3i][x-1+3i] [(x-1)-(3i)][(x-1)+(3i)] do you see the difference of squares (a-b)(a+b)?

OpenStudy (precal):

[(x-1)^2-9i^2] x^2-2x+1-9i^2 note i^2 is -1 x^2-2x+1-9(-1) x^2-2x+1+9 x^2-2x+10 ok so now multiply this to your other two

OpenStudy (precal):

you can always verify that this is true by solving for x using the quadratic formula

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (precal):

yw